Grade 12th passPhysical Chemistry

Vantshoff factor of hg2cl2 in its solution will be if 80./.ionized in the solution ?

Mamta thakur
8 Years agoGrade 12th pass

1 Answer

Hemant
8 Years ago
dissociation constant =i-1/n-10.8=i-1/4-1i=3.4n=4 because hg2cl2 wipl dissociate into 4 ionsi is the vonthoff factor